# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 8.3. Approximation by coarse-grained model
Indeed, the second moment can be computed as in Subsection 8.2. For the fourth moment, similar to Lemma 6.2, if we already compute the fourth moment of $\Theta(\vec{\mathrm{i}})$ for $|\vec{\mathrm{i}}|\leq 2$, then for $|\vec{\mathrm{i}}|\geq 3$, we can decouple $\Theta(\mathrm{i},\mathrm{i})$ and $\Theta(\mathrm{i}^{\prime},\mathrm{i}^{\prime})$ and the result follows immediately. Hence, it only remains to compute the case $|\vec{\mathrm{i}}|\leq 2$. Without loss of generality, we only need to treat the case $|\vec{\mathrm{i}}|=1$.
We have that by Gaussian property and Wick formula, ${\mathbbm{E}}[(\Theta(i,i))^{4}]$ is bounded above by
(8.24) $\begin{split}\frac{3\delta^{4}}{\varepsilon^{2}}\sum\limits_{k=2}^{\infty}\beta_{\delta}^{4k}&\Bigg{(}\iint_{{\mathbbm{R}}^{2}}\rho(x_{1})\rho(y_{1}){\mathrm{d}}x_{1}{\mathrm{d}}y_{1}\idotsint\limits_{t_{1}<\cdots<t_{k}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i})\atop x_{2},y_{2},\cdots,x_{k},y_{k}\in{\mathbbm{R}}}\prod\limits_{i=2}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\times\\ &\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}\Bigg{)}^{2}+\frac{\delta^{4}}{\varepsilon^{2}}\Bigg{(}\sum\limits_{m=2}^{\infty}\beta_{\delta}^{2m}\iint_{{\mathbbm{R}}^{2}}\rho(x_{1})\rho(y_{1}){\mathrm{d}}x_{1}{\mathrm{d}}y_{1}\idotsint\limits_{t_{1}<\cdots<t_{m}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i})\atop x_{2},y_{2},\cdots,x_{m},y_{m}\in{\mathbbm{R}}}\\ &\prod\limits_{i=2}^{m}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}\Bigg{)}^{2}.\end{split}$
As $\delta\to 0$, the second term above converges to
(8.25) $\begin{split}\frac{4\pi}{\varepsilon^{2}}\Big{(}\iint_{(\mathrm{i}-1)\varepsilon<s<t<\mathrm{i}\varepsilon}G_{\vartheta}(t-s){\mathrm{d}}s{\mathrm{d}}t\Big{)}^{2}\leq\frac{C}{(\log\frac{1}{\varepsilon})^{2}}.\end{split}$
For the first term, by (8.6)-(8.9), as $\delta\to 0$, it equals to
(8.26) $\begin{split}\lim\limits_{\delta\to 0}&\frac{3}{\varepsilon^{2}}\sum\limits_{k=2}^{\infty}\lambda_{\delta}^{2k}\bigg{(}\int_{(\mathrm{i}-1)\varepsilon}^{\mathrm{i}\varepsilon}{\mathbf{P}}\Big{(}{\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{k}^{\delta}\leq\delta^{-2}(1-t)\Big{)}{\mathrm{d}}t\bigg{)}^{2}\\ &=\frac{3}{\varepsilon^{2}}\int_{0}^{+\infty}e^{2\vartheta u}\bigg{(}\int_{(\mathrm{i}-1)\varepsilon}^{\mathrm{i}\varepsilon}{\mathrm{d}}t\int_{0}^{1-t}\alpha_{u}(s){\mathrm{d}}s\bigg{)}^{2}{\mathrm{d}}u.\end{split}$
Note that by (2.25), uniformly for $u\geq 0$,
$e^{\vartheta u}\int_{(\mathrm{i}-1)\varepsilon}^{\mathrm{i}\varepsilon}{\mathrm{d}}t\int_{0}^{1-t}\alpha_{u}(s){\mathrm{d}}s\leq\varepsilon e^{\vartheta u}\int_{0}^{1}\frac{e^{-\gamma u}us^{u-1}}{\Gamma(u+1)}{\mathrm{d}}s=\varepsilon\frac{e^{(\vartheta-\gamma)u}}{\Gamma(u+1)}\leq C\varepsilon.$
Hence, (8.26) is bounded above by
$\frac{C}{\varepsilon}\int_{0}^{+\infty}e^{\vartheta u}{\mathrm{d}}u\int_{(\mathrm{i}-1)\varepsilon}^{\mathrm{i}\varepsilon}{\mathrm{d}}t\int_{0}^{1-t}\alpha_{u}(s){\mathrm{d}}s=\frac{C}{\varepsilon}\iint_{(\mathrm{i}-1)\varepsilon<s<t<\mathrm{i}\varepsilon}G_{\vartheta}(t-s){\mathrm{d}}s{\mathrm{d}}t\leq\frac{C}{\log\frac{1}{\varepsilon}},$
which completes the proof.
Acknowledgment
R. Wei is supported by XJTLU Research Development Fund (No. RDF-23-01-024). J. Yu is supported by National Key R&D Program of China (No. 2021YFA1002700) and NSFC (No. 12101238). We thank Francesco Caravenna for very helpful discussion and for pointing out the connection between disordered pinning models and stochastic Volterra equations. We also thank Quentin Berger, Rongfeng Sun and Nikolaos Zygouras for reading the first version of the paper and providing invaluable suggestions.
@@ Appendix A Enhanced Lindeberg Principle
To make this paper more self-contained, we cite [CSZ23, Appendix A] here.
@@
(A.1) $W_{t,u}^{k,l}:=u\sqrt{t}X^{A_{kl}}+\sqrt{t}A^{A_{kl}^{c}}+\sqrt{1-t}Z,$
@@
(A.2) $W_{s,t,u}^{k,l}:=su\sqrt{t}X^{A_{k}}+u\sqrt{t}X^{A_{kl}\backslash A_{k}}+\sqrt{t}X^{A_{kl}^{c}}+\sqrt{1-t}Z.$
We will apply following lemmas in proving our main results, which are [CSZ23, Lemma A.2-A.4]. Note that we use $\partial_{k}$ to denote the partial derivative with respect to $X_{k}$.
@@
(A.3) $I_{1}:=\frac{1}{2}\iiint_{[0,1]^{3}}$ $\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}{\mathbbm{E}}\Big{[}X_{l}X_{k}X_{m}(s\mathbbm{1}_{\{m\in A_{K}\}})+\mathbbm{1}_{\{m\in A_{kl}\backslash A_{k}\}}\sqrt{t}\partial^{3}_{klm}h\big{(}W_{s,t,u}^{k,l}\big{)}\Big{]}{\mathrm{d}}s{\mathrm{d}}t{\mathrm{d}}u,$
(A.4) $I_{2}$ $:=-\frac{1}{2}\iint_{[0,1]^{2}}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}\sigma_{kl}{\mathbbm{E}}\Big{[}X_{m}\sqrt{t}\partial^{3}_{klm}h\big{(}W_{t,u}^{k,l}\big{)}\Big{]}{\mathrm{d}}t{\mathrm{d}}u.$
@@
(A.5) ${\mathbbm{E}}[h(Z)]-{\mathbbm{E}}[h(\widetilde{Z})]=I_{3}:=\frac{1}{2}\sum_{k,l\in{\mathbbm{T}}}(\widetilde{\sigma}_{kl}-\sigma_{kl})\int_{0}^{1}{\mathbbm{E}}\big{[}\partial^{2}_{kl}h(W_{t})\big{]}{\mathrm{d}}t,$
@@
$\Lambda(X)=\sum\limits_{I\subset{\mathbbm{T}}}c_{I}\prod_{i\in I}x_{i}$
@@
(A.6) $\forall~{}k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl},\quad\partial_{km}^{2}\Lambda=\partial_{lm}^{2}\Lambda=\partial_{kl}^{2}\Lambda=0.$
@@
(A.7) $|I_{1}|\leq\frac{1}{2}\|$ $f^{\prime\prime\prime}\|_{\infty}\sup_{k\in{\mathbbm{T}}}{\mathbbm{E}}\big{[}|X_{k}|^{3}\big{]}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}$
@@
(A.8) $|I_{2}|\leq\frac{1}{2}\|$ $f^{\prime\prime\prime}\|_{\infty}\sup_{k\in{\mathbbm{T}}}{\mathbbm{E}}\big{[}|X_{k}|^{3}\big{]}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}$
@@
(A.9) $|I_{3}|\leq\frac{1}{2}\|$ $f^{\prime\prime}\|_{\infty}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k}}(\widetilde{\sigma}_{kl}-\sigma_{kl})\sup_{t}{\mathbbm{E}}\Big{[}\big{|}\partial_{k}\Lambda(W_{t})\big{|}^{2}\Big{]}^{\frac{1}{2}}\sup_{t}{\mathbbm{E}}\Big{[}\big{|}\partial_{l}\Lambda(W_{t})\big{|}^{2}\Big{]}^{\frac{1}{2}}.$
@@
where $k$ is a finite positive integer and $f:{\mathbbm{R}}^{k}\to k$ is a bounded function with bounded derivatives up to third order. Similar bounds as (A.7)-(A.9) still hold. For (A.7) and (A.8), we need to replace $\|f^{\prime\prime\prime}\|_{\infty}$ by $\max_{1\leq i,j,l\leq k}\|\partial_{i,j,l}f\|_{\infty}$, replace the three occurrences of $\Lambda$ by $\Lambda_{i},\Lambda_{j}$ and $\Lambda_{l}$, and sum over $1\leq i,j,l\leq k$. For (A.9), we need to replace $\|f^{\prime\prime}\|_{\infty}$ by $\max_{1\leq i,j\leq k}$, replace the two occurrences of $\Lambda$ by $\Lambda_{i}$ and $\Lambda_{j}$, and sum over $1\leq i,j\leq k$.

### v3

@@ 8.3 Approximation by the coarse-grained model
Indeed, the second moment can be computed as in Subsection 8.2. For the fourth moment, similar to Lemma 6.2, if we already compute the fourth moment of $\Theta(\vec{\mathrm{i}})$ for $|\vec{\mathrm{i}}|\leq 2$, then for $|\vec{\mathrm{i}}|\geq 3$, we can decouple $\Theta(\mathrm{i},\mathrm{i})$ and $\Theta(\mathrm{i}^{\prime},\mathrm{i}^{\prime})$ and the result follows immediately. Hence, it only remains to compute the case $|\vec{\mathrm{i}}|\leq 2$. Without loss of generality, we only need to treat the case $|\vec{\mathrm{i}}|=1$. We explain the methodology below and omit tedious computations.
First note that by the computations in Sections 8.1-8.2, we can restrict the number of transition kernels $g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i})$ in (8.22) by $(\log\delta^{-1})^{2}$ for $\delta$ small enough. Consequently, since $\Theta(\vec{\mathrm{i}})$ is a block in (8.22), we can also apply these restrictions to $\Theta(\vec{\mathrm{i}})$. When expanding $\Theta(\vec{\mathrm{i}})$, we obtain a summation with at most $(\log\delta^{-1})^{8}$ terms, where each summand is the product of four multiple stochastic integrals with orders $k_{i},i=1,2,3,4$. We merge each summand to a $\sum_{i=1}^{4}k_{i}$-multiple stochastic integral. Then by the property of ${\mathrm{d}}W_{t}$, if a summand contains a ${\mathrm{d}}W_{t}$, $({\mathrm{d}}W_{t})^{3}$ or $({\mathrm{d}}W_{t})^{4}$ term, then it has 0 expectation (we can switch the expectation and summation since there are only finite many terms). Hence, by comparing to the discrete case, now we have four Brownian motions and we should record the time when exactly two of them visit the support of the mollifier $\rho$ (we may call it “collision”).
Next, we can rearrange the summation according to the recording time above, just as we have done in Section 4 and Section 6. Also note that now we only have counterparts of function $U$ in (4.11) or type $U$ groups in Section 6, since other cases therein refer to a collision for at least three Brownian motions. Then the bound similar to Lemma 6.2 can be obtained by the frameworks in Section 4 and Section 6. The details for computing the collision for two Brownian motions have been given in Section 8.2 (recall that this collision is closely related to the Dickman subordinator).
@@ Appendix A Enhanced Lindeberg Principle
To make this paper more self-contained, we cite [15, Appendix A] here.
@@
$W_{t,u}^{k,l}:=u\sqrt{t}X^{A_{kl}}+\sqrt{t}A^{A_{kl}^{c}}+\sqrt{1-t}Z,$ (A.1)
@@
$W_{s,t,u}^{k,l}:=su\sqrt{t}X^{A_{k}}+u\sqrt{t}X^{A_{kl}\backslash A_{k}}+\sqrt{t}X^{A_{kl}^{c}}+\sqrt{1-t}Z.$ (A.2)
We will apply following lemmas in proving our main results, which are [15, Lemma A.2-A.4]. Note that we use $\partial_{k}$ to denote the partial derivative w.r.t. $X_{k}$.
@@
$I_{1}:=$ $\frac{1}{2}\iiint_{[0,1]^{3}}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}$ (A.3)
${\mathbbm{E}}\Big[X_{l}X_{k}X_{m}(s\mathbbm{1}_{\{m\in A_{K}\}})+\mathbbm{1}_{\{m\in A_{kl}\backslash A_{k}\}}\sqrt{t}\partial^{3}_{klm}h\big(W_{s,t,u}^{k,l}\big)\Big]{\mathrm{d}}s{\mathrm{d}}t{\mathrm{d}}u,$
$I_{2}:=$ $-\frac{1}{2}\iint_{[0,1]^{2}}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}\sigma_{kl}{\mathbbm{E}}\Big[X_{m}\sqrt{t}\partial^{3}_{klm}h\big(W_{t,u}^{k,l}\big)\Big]{\mathrm{d}}t{\mathrm{d}}u.$ (A.4)
@@
${\mathbbm{E}}[h(Z)]-{\mathbbm{E}}[h(\widetilde{Z})]=I_{3}:=\frac{1}{2}\sum_{k,l\in{\mathbbm{T}}}(\widetilde{\sigma}_{kl}-\sigma_{kl})\int_{0}^{1}{\mathbbm{E}}\big[\partial^{2}_{kl}h(W_{t})\big]{\mathrm{d}}t,$ (A.5)
@@
$\Lambda(X)=\sum\limits_{I\subset{\mathbbm{T}}}c_{I}\prod_{i\in I}x_{i},$
@@
$\forall~k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl},\quad\partial_{km}^{2}\Lambda=\partial_{lm}^{2}\Lambda=\partial_{kl}^{2}\Lambda=0.$ (A.6)
@@
$|I_{1}|\leq\frac{1}{2}\|$ $f^{\prime\prime\prime}\|_{\infty}\sup_{k\in{\mathbbm{T}}}{\mathbbm{E}}\big[|X_{k}|^{3}\big]\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}$ (A.7)
@@
$|I_{2}|\leq\frac{1}{2}\|$ $f^{\prime\prime\prime}\|_{\infty}\sup_{k\in{\mathbbm{T}}}{\mathbbm{E}}\big[|X_{k}|^{3}\big]\sum\limits_{k\in{\mathbbm{T}},l\in A_{k},m\in A_{kl}}$ (A.8)
@@
$|I_{3}|\leq\frac{1}{2}\|$ $f^{\prime\prime}\|_{\infty}\sum\limits_{k\in{\mathbbm{T}},l\in A_{k}}(\widetilde{\sigma}_{kl}-\sigma_{kl})\sup_{t}{\mathbbm{E}}\Big[\big|\partial_{k}\Lambda(W_{t})\big|^{2}\Big]^{\frac{1}{2}}\sup_{t}{\mathbbm{E}}\Big[\big|\partial_{l}\Lambda(W_{t})\big|^{2}\Big]^{\frac{1}{2}}.$ (A.9)
@@
where $k$ is a finite positive integer and $f:{\mathbbm{R}}^{k}\to{\mathbbm{R}}$ is a bounded function with bounded derivatives up to third order. Similar bounds as (A.7)-(A.9) still hold. For (A.7) and (A.8), we need to replace $\|f^{\prime\prime\prime}\|_{\infty}$ by $\max_{1\leq i,j,l\leq k}\|\partial_{ijl}^{3}f\|_{\infty}$, replace the three occurrences of $\Lambda$ by $\Lambda_{i},\Lambda_{j}$ and $\Lambda_{l}$, and sum over $1\leq i,j,l\leq k$. For (A.9), we need to replace $\|f^{\prime\prime}\|_{\infty}$ by $\max_{1\leq i,j\leq k}\|\partial_{ij}^{2}f\|_{\infty}$, replace the two occurrences of $\Lambda$ by $\Lambda_{i}$ and $\Lambda_{j}$, and sum over $1\leq i,j\leq k$.
Appendix B Errors of Local Limit Theorem
The following result is well-known (see [27, Theorem 4.5.3]).
Theorem B.1.
Let $S$ be an irreducible and aperiodic random walk specified in Theorem 1.1. Then for any $\kappa>0$ in (1.10),
$\sup_{k\in{\mathbbm{N}}}\Big|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\Big(\frac{k}{\sqrt{n}}\Big)\Big|=O(n^{-\frac{\kappa\wedge 1}{2}}),$ (B.1)
where $\varphi$ is the density of the standard Gaussian distribution.
Conversely, if a random walk $S$ satisfies (B.1) for $0<\kappa<1$, then ${\mathbf{E}}[|S_{1}|^{2}\mathbbm{1}_{\{|S_{1}|>k\}}]=O(k^{-\kappa})$ and $S$ has span $1$.
Then we can prove the following refined local limit theorem, which to the best of our knowledge, has not appeared in literature.
Theorem B.2.
Let $S$ be as specified in Theorem B.1. Then for any $\kappa>0$ in (1.10), there exist constants $\gamma,C>0$, such that uniformly in $n\in{\mathbbm{N}}$ and $k\in{\mathbbm{Z}}$,
$\Big|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\Big(\frac{k}{\sqrt{n}}\Big)\Big|\leq\frac{C}{1+|k|^{1+\gamma}}.$ (B.2)
Proof.
First, we have that for any $0<\theta<2$,
$\lim\limits_{n\to\infty}\sup\limits_{k\in{\mathbbm{Z}}}\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\Big|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\Big(\frac{k}{\sqrt{n}}\Big)\Big|=0,$ (B.3)
which is a special case of [30, Theorem 1]. We present a proof of (B.3) here for the completeness. Denote $K_{n}(A):=\{k\in{\mathbbm{Z}}:k\geq A\sqrt{n}\}$ for $A>0$ and
$\Delta_{n,k}^{\theta}:=\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\Big|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\Big(\frac{k}{\sqrt{n}}\Big)\Big|.$
Then
$\begin{split}\sup\limits_{k\in{\mathbbm{Z}}}\Delta_{n,k}^{\theta}\leq&\sup_{k\in K_{n}(A)}\Delta_{n,k}^{\theta}+\sup_{k\in K_{n}(A)^{c}}\Delta_{n,k}^{\theta}\\ \leq&\sup_{k\in K_{n}(A)}\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\sqrt{n}{\mathbf{P}}(S_{n}=k)+\sup_{k\in K_{n}(A)}\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\varphi\Big(\frac{k}{\sqrt{n}}\Big)\\ &+A^{\theta}\sup_{k\in K_{n}(A)^{c}}\Big|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\Big(\frac{k}{\sqrt{n}}\Big)\Big|.\end{split}$
For any $\varepsilon>0$, by choosing $A$ large, the second term is smaller than $\varepsilon$ for all $n$. For fixed $A$, the last term is smaller than $\varepsilon$ for large enough $n$ by the classic local limit theorem. We are left to treat the first term above. We only need to show that for any $0<\theta<2$,
$\sup_{n\in{\mathbbm{N}}}\sup_{k\in{\mathbbm{Z}}}\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\sqrt{n}{\mathbf{P}}(S_{n}=k)<\infty,$ (B.4)
since we can apply (B.4) to some $\theta^{\prime}\in(\theta,2)$ and note that $\sup_{k\in K_{n}(A)}|k/\sqrt{n}|^{\theta-\theta^{\prime}}\leq A^{\theta-\theta^{\prime}}$ is smaller than $\varepsilon$ by choosing $A$ large.
We now prove (B.4). By [27, Lemma 5.2.2], for any $0<\theta<2$,
$\sup\limits_{n\in{\mathbbm{Z}}}{\mathbf{E}}\bigg[\bigg|\frac{S_{n}}{\sqrt{n}}\bigg|^{\theta}\bigg]=\sup\limits_{n\in{\mathbbm{N}}}\sum\limits_{k\in{\mathbbm{Z}}}\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}{\mathbf{P}}(S_{n}=k)<\infty.$ (B.5)
Write
$\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\sqrt{n}{\mathbf{P}}(S_{n}=k)=\sum\limits_{\ell\in{\mathbbm{Z}}}\bigg|\frac{k}{\sqrt{n}}\bigg|^{\theta}\sqrt{n}{\mathbf{P}}(S_{\frac{n}{2}}=\ell){\mathbf{P}}(S_{\frac{n}{2}}=k-\ell).$ (B.6)
For any $\theta>0$, we have $|k|^{\theta}\leq 2^{\theta}\big(|\ell|^{\theta}+|k-\ell|^{\theta}\big)$. To see this, note that for any $x,y\in{\mathbbm{R}}$, $|x+y|^{\theta}\leq 2^{\theta}(\max\{|x|,|y|\})^{\theta}\leq 2^{\theta}(|x|^{\theta}+|y|^{\theta})$. Hence, by the classic local limit theorem, $\sup_{n}\sup_{k}\sqrt{n}{\mathbf{P}}(S_{n/2}=k)\leq C$, and thus (B.6) is bounded by $2^{\theta}$ times
$\begin{split}&\sum\limits_{\ell\in{\mathbbm{Z}}}\bigg(\bigg|\frac{\ell}{\sqrt{n}}\bigg|^{\theta}+\bigg|\frac{k-\ell}{\sqrt{n}}\bigg|^{\theta}\bigg)\sqrt{n}{\mathbf{P}}(S_{\frac{n}{2}}=\ell){\mathbf{P}}(S_{\frac{n}{2}}=k-\ell)\\ \leq&C\sum\limits_{\ell\in{\mathbbm{Z}}}\bigg|\frac{\ell}{\sqrt{n}}\bigg|^{\theta}{\mathbf{P}}(S_{\frac{n}{2}}=\ell)+C\sum\limits_{\ell\in{\mathbbm{Z}}}\bigg|\frac{k-\ell}{\sqrt{n}}\bigg|^{\theta}{\mathbf{P}}(S_{\frac{n}{2}}=k-\ell),\end{split}$
and then (B.4) is proved by (B.5).
With (B.1) and (B.3) at hand, we now prove (B.2). It suffices to treat small $\kappa>0$. We show that $\gamma=\frac{\kappa^{2}}{6}$ is valid for (B.2). We consider two cases: (1) $|k|\leq n^{\frac{1}{2}+\frac{\kappa}{4}}$ and (2) $|k|\geq n^{\frac{1}{2}+\frac{\kappa}{4}}$.
For case (1), note that
$\bigg|\frac{k}{\sqrt{n}}\bigg|^{1+\gamma}\leq n^{\frac{\kappa}{4}(1+\frac{\kappa^{2}}{6})}\ll n^{\frac{\kappa}{2}},$
and then (B.2) follows from (B.1).
For case (2), apply (B.3) to $\theta=\kappa$,
$\begin{split}o(1)&=\sup\limits_{|k|\geq n^{\frac{1}{2}+\frac{\kappa}{4}}}\bigg|\frac{k}{\sqrt{n}}\bigg|^{1+\kappa}\bigg|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\bigg(\frac{k}{\sqrt{n}}\bigg)\bigg|\\ &=\sup\limits_{|k|\geq n^{\frac{1}{2}+\frac{\kappa}{4}}}\frac{|k|^{1+\frac{\kappa^{2}}{6}}}{\sqrt{n}}\frac{|k|^{\kappa-\frac{\kappa^{2}}{6}}}{\sqrt{n}^{\kappa}}\bigg|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\bigg(\frac{k}{\sqrt{n}}\bigg)\bigg|.\end{split}$
Note that in case (2),
$\frac{|k|^{\kappa-\frac{\kappa^{2}}{6}}}{\sqrt{n}^{\kappa}}\geq n^{\frac{\kappa^{2}}{3}-\frac{\kappa^{3}}{24}}\longrightarrow\infty,\quad\text{as}~n\to\infty.$
Hence, we must have
$o(1)=\sup\limits_{|k|\geq n^{\frac{1}{2}+\frac{\kappa}{4}}}\frac{|k|^{1+\frac{\kappa^{2}}{6}}}{\sqrt{n}}\bigg|\sqrt{n}{\mathbf{P}}(S_{n}=k)-\varphi\bigg(\frac{k}{\sqrt{n}}\bigg)\bigg|,$
which completes the proof. ∎
{acks}
[Acknowledgments] We thank Francesco Caravenna for very helpful discussion and for pointing out the connection between disordered pinning models and stochastic Volterra equations. We also thank Quentin Berger, Rongfeng Sun and Nikolaos Zygouras for reading the first version of the paper and providing invaluable suggestions. Finally, we thank the anonymous referee who helps us significantly improve the quality of this paper. J. Yu acknowledges the support of NYU-ECNU Institute of Mathematical Sciences at NYU Shanghai.
{funding}
R. Wei is supported by XJTLU Research Development Fund (No. RDF-23-01-024) and NSFC (No. 12401170). J. Yu is supported by National Key R&D Program of China (No. 2021YFA1002700) and NSFC (No. 12101238 and No. 12271010).